# BINDy – Bayesian identification of nonlinear dynamics with reversible-jump Markov-chain Monte-Carlo

### v2

@@ 3 Bayesian identification of nonlinear dynamics
$\Sigma_{m}=\sigma^{2}\left[\Theta_{m}^{\mkern-1.5mu\mathsf{T}}\Theta_{m}+\Sigma_{m}^{(0)-1}\right]^{-1}{\color[rgb]{0,0,0},}$ (19)
@@
$\begin{split}p(m|\mathcal{D},\sigma^{2})&\propto p(m)p(\mathcal{D},\sigma^{2}|m,\Xi){\color[rgb]{0,0,0},}\\ &\propto p(m)\sigma^{-k}|\Sigma_{m}|^{\frac{1}{2}}\exp\left(\frac{1}{2}\mu_{m}^{\mkern-1.5mu\mathsf{T}}\Sigma^{-1}_{m}\mu_{m}\right){\color[rgb]{0,0,0}.}\end{split}$ (21)
Thus, the overall acceptance probability of a move from $\Xi$ to $\Xi^{\prime}$ can be written,
$\alpha(\Xi\rightarrow\Xi^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{p(m^{\prime})}{p(m)}}_{\text{Model prior}}\underbrace{\frac{J(m^{\prime}|m)}{J(m|m^{\prime})}}_{\text{Jump ratio}}\underbrace{\frac{\sigma^{-k^{\prime}}|\Sigma_{m^{\prime}}|^{\frac{1}{2}}\exp\left(\frac{1}{2}\mu_{m^{\prime}}^{\mkern-1.5mu\mathsf{T}}\Sigma^{-1}_{m^{\prime}}\mu_{m^{\prime}}\right)}{\sigma^{-k}|\Sigma_{m}|^{\frac{1}{2}}\exp\left(\frac{1}{2}\mu_{m}^{\mkern-1.5mu\mathsf{T}}\Sigma^{-1}_{m}\mu_{m}\right)}}_{\begin{subarray}{c}\text{Model posterior ratio}\end{subarray}}\Big{\}}{\color[rgb]{0,0,0}.}$ (22)
The acceptance probability in (22) can now be readily used to sample from the required posterior. In practice, the noise variance in (3) will be unknown. This problem can be simply addressed by including a Gibbs move for the noise variance given the model and the parameters, which are available following a reversible-jump move. For example an inverse-gamma prior, conjugate to the Gaussian likelihood can be employed. The Gibbs step in this case is,
@@ 4.3 Lorenz attractor dynamics
The result of the UQ for the Lorenz equations is compared in Figure 5 for BINDy (with and without the regularising prior) and E-SINDy. It can be seen in the figure that the unregularised BINDy assigns some posterior probability to several spurious terms, particularly in the third equation. In contrast, Both E-SINDy and the regularised BINDy assign high confidence to terms in the Lorenz equations. This is not an unexpected result. The nature of the Lorenz equations, and the use of the polynomial library results in many, highly-correlated basis functions in $\Theta(x)$. This, plus the effect of noise permit many incorrect (but low error) parametrisations of the observed dynamics. Both E-SINDy and regularised BINDy are able to overcome this difficulty by explicitly promoting model parsimony. Because the unregularised BINDy enforces no such cognitive bias, it is unsurprising that the identified posterior contains many terms. Summary statistics of the three identified posteriors are provided in Table 1.
As before, sampled model trajectories (integrated forward in time) are plotted for all three methods in Figure 6. The chaotic nature of the Lorenz equations makes tracking the true dynamics highly challenging. The well-known ‘lobe-switching’ behavior of the dynamics means that small errors can quickly cause significant deviations as the dynamics become governed by a different local attractor. This effect is clearly present in the identified distributions, which quickly become multimodal after a few oscillations. Qualitatively, it is observed that the regularised BINDy samples remain in step with the true dynamics for the longest, with E-SINDy the first to diverge. This divergence can be seen almost immediately where some orange lines deviate quickly from the observed data. Indeed, this is supported by the numerical results in Table 2 which show lower error statistics for the regularised BINDy approach. In the extrapolation region, the distributions have become very diffuse, in keeping with what might be expected for uncertain forecasting of a chaotic attractor such as the Lorenz equations.
@@ 5 Discussion
In the first two case studies, strongly parsimonious models were identified by the proposed methodology. Counterintuitively, these models were assigned high posterior probability despite a flat prior over the model space. It is interesting to consider how this can be the case, for which the authors offer two explanations. On examination of the acceptance ratio in (22), one finds that the ratio is proportional to $\sigma^{-k^{\prime}}/\sigma^{-k}$. Intuitively, this ratio has a slight regularising effect that penalises larger models that have the same fit to the data. Another important consideration is the correlations between the columns of $\Theta(x)$ and the state derivatives $\dot{x}$. These correlations depend on the data, the system under study and chosen library of basis functions. Many, highly correlated terms can complicate the identification procedure with many models able to produce predictions close to the observed data. In contrast, in the case where only a few terms have high correlation, the identification of models with few terms is simplified. This effect could also explain the parsimonious models in the first two case studies.
@@
As has been demonstrated in the third case-study of this work, the choice of prior over models $p(m)$ may be critical in enforcing parsimony. Here, a strongly regularising geometric distribution parameterised by the number of terms in the model has been applied successfully. However, vastly many more parametrisations are possible. It is expected that prior selection in practice will be driven by expert domain knowledge, wherever possible. For example, one could set the prior over each term directly, assigning high prior probability to terms that are expected to be in the model and low probabilities otherwise. Alternatively one could promote parsimony in some types of library functions while permitting dense models in other types. The flexible nature of the Bayesian formulation presented in this work, allows modellers to encode both domain knowledge and cognitive biases a priori. An interesting avenue for further investigation is hierarchical prior structures whereby the model-space prior is permitted to be conditional on the values its parametrisation.
@@ Appendix C Robustness of BINDy
Nevertheless, it is instructive to consider the effect of low-data and high-noise regimes, particularly in the context of the sparsity-promoting priors in the model space that were applied in the Lorenz case study example. In Figures 10 and 11, the posterior term inclusion probabilities are plotted. BINDy is applied to the Lorenz case-study example both with (green) and without (blue) the regularising prior. In Figure 10 the length of the observed data is held constant at 10s (or 1000 observations) and the variance of the added noise is varied from 1% to 15%. As expected, the posterior distribution resulting from BINDy with a flat model prior becomes very diffuse as the noise level increases, indicating that the added noise conveys significant uncertainty as to which terms should be included in the model. At low noise levels, the flat model prior correctly assigns high probability to the correct model. Also in Figure 10, the effect of the regularising is seen to have a dramatic effect on the posterior term inclusion probabilities. Even at high noise levels, the correct terms are assigned high probabilities with some spurious terms appearing at the highest noise levels. Once again at low noise levels, the regularised prior assigns almost all probability mass to the correct model terms.
Figure 10: Posterior term inclusion probabilities at a fixed data length for various levels of added Gaussian noise. Estimated by BINDy with a flat model prior (green) and a regularising prior (blue). Subplot axes are identical to those of Figure 5. Solid bars indicate terms truly in the data generating model, while hollow bars indicate spurious terms.
In Figure 11, the effect of decreasing amounts of available data is investigated. The level of added noise is held constant at 2.5% RMS and the length of the observed dynamics available for inference is varied between 1.0 seconds (corresponding to only 100 datapoints) up to 10 seconds (corresponding to 1000 datapoints). Once again, both BINDy with a flat and regularising model prior are considered. As can be seen in the Figure, the posterior term inclusion probabilities estimated by BINDy without the regularising prior are very diffuse when only a few data are available, corresponding to the broad likelihood. As the amount of data available increases and the likelihood contracts, it is observed that the posterior distributions contract around the true model parameters, as might be expected of a Bayesian approach such as this. The posterior term inclusion probabilities estimated by BINDy with the regularising prior show remarkable robustness to the length of available data and assigning almost all posterior probability mass to the correct terms in most cases.
Figure 11: Posterior term inclusion probabilities at a fixed noise level for various amounts of observed dynamics. Estimated by BINDy with a flat model prior (green) and a regularising prior (blue). Subplot axes are identical to those of Figure 5. Solid bars indicate terms truly in the data generating model, while hollow bars indicate spurious terms.
Figure 12 depicts the robustness of the regularising prior to the effect of data length and noise in a heatmap format. Each element of the heatmap considers the probability of the true data-generating model under the posterior estimated by BINDy with the regularising prior. As can be seen in the figure, the probability approaches unity for low noise and long sequences.
Figure 12: Heatmap depicting the posterior probability of the true model with the regularising prior for different noise levels and data lengths.

### v3

@@ 3 Bayesian identification of nonlinear dynamics
$\Sigma_{m}=\left[\frac{1}{\sigma^{2}}\Theta_{m}^{\mkern-1.5mu\mathsf{T}}\Theta_{m}+\Sigma_{m}^{(0)-1}\right]^{-1}{\color[rgb]{0,0,0},}$ (19)
@@
$\begin{split}p(m|\mathcal{D},\sigma^{2})&\propto p(m)p(\mathcal{D}|m,\sigma^{2}){\color[rgb]{0,0,0},}\\ &=p(m)(2\pi\sigma^{2})^{\frac{N}{2}}|\Sigma^{(0)}_{m}|^{-\frac{1}{2}}|\Sigma_{m}|^{\frac{1}{2}}\exp\left(-\frac{1}{2}\left[\sigma^{-2}\mathcal{D}^{\mkern-1.5mu\mathsf{T}}\mathcal{D}-\mu_{m}^{\mkern-1.5mu\mathsf{T}}\Sigma^{-1}_{m}\mu_{m}+[\mu^{(0)}_{m}]^{\mkern-1.5mu\mathsf{T}}[\Sigma_{m}^{(0)}]^{-1}\mu^{(0)}_{m}\right]\right){\color[rgb]{0,0,0}.}\end{split}$ (21)
Now assuming a zero-mean prior over the model parameters $\Xi$ [^3] Note that in the case of non-zero mean parameter priors, an additional exponentiated quadratic form will appear in the numerator and denominator of the acceptance ratio., the overall acceptance probability of a move from $\Xi$ to $\Xi^{\prime}$ can be written,
$\alpha(\Xi\rightarrow\Xi^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{p(m^{\prime})}{p(m)}}_{\text{Model prior}}\underbrace{\frac{J(m^{\prime}|m)}{J(m|m^{\prime})}}_{\text{Jump ratio}}\underbrace{\frac{|\Sigma^{(0)}_{m^{\prime}}|^{-\frac{1}{2}}|\Sigma_{m^{\prime}}|^{\frac{1}{2}}\exp\left(\frac{1}{2}\mu_{m^{\prime}}^{\mkern-1.5mu\mathsf{T}}\Sigma^{-1}_{m^{\prime}}\mu_{m^{\prime}}\right)}{|\Sigma^{(0)}_{m}|^{-\frac{1}{2}}|\Sigma_{m}|^{\frac{1}{2}}\exp\left(\frac{1}{2}\mu_{m}^{\mkern-1.5mu\mathsf{T}}\Sigma^{-1}_{m}\mu_{m}\right)}}_{\begin{subarray}{c}\text{Model posterior ratio}\end{subarray}}\Big{\}},$ (22)
where several terms that do not depend on $m$ have cancelled out in the model posterior ratio. The acceptance probability in (22) can now be readily used to sample from the required posterior. In practice, the noise variance in (3) will be unknown. This problem can be simply addressed by including a Gibbs move for the noise variance given the model and the parameters, which are available following a reversible-jump move. For example an inverse-gamma prior, conjugate to the Gaussian likelihood can be employed. The Gibbs step in this case is,
@@ 4.3 Lorenz attractor dynamics
The result of the UQ for the Lorenz equations is compared in Figure 5 for BINDy (with and without the regularising prior) and E-SINDy. It can be seen in the figure that all three approaches have been highly successful in assigning high probability to the terms in the data-generating model. As before, sampled model trajectories (integrated forward in time) are plotted for all three methods in Figure 6. The chaotic nature of the Lorenz equations makes tracking the true dynamics highly challenging. The well-known ‘lobe-switching’ behavior of the dynamics means that small errors can quickly cause significant deviations as the dynamics become governed by a different local attractor. This effect is clearly present in the identified distributions, which quickly become multimodal after a few oscillations. Qualitatively, it is observed that both BINDy and regularised BINDy samples remain in step with the true dynamics for the longest, with E-SINDy the first to diverge. This divergence can be seen almost immediately where some orange lines deviate quickly from the observed data. Indeed, this is supported by the numerical results in Table 2 which show lower error statistics for the BINDy and regularised BINDy approach. In the extrapolation region, the distributions have become very diffuse, in keeping with what might be expected for uncertain forecasting of a chaotic attractor such as the Lorenz equations.
@@ 5 Discussion
In all three case studies, strongly parsimonious models were identified by the proposed methodology. Counterintuitively, these models were assigned high posterior probability despite a flat prior over the model space. It is interesting to consider how this can be the case, for which the authors offer two explanations. On examination of the acceptance ratio in (22), one finds that the ratio is proportional to $|\Sigma^{(0)}_{m^{\prime}}|^{-\frac{1}{2}}/|\Sigma^{(0)}_{m}|^{-\frac{1}{2}}$. Intuitively, this ratio has a regularising effect that penalises larger models that have the same fit to the data due to the differing dimensions of the prior covariance. This effect scales with the parameter covariance meaning that wide priors (like those considered in the case studies here) have a stronger regularising effect. Another important consideration is the correlations between the columns of $\Theta(x)$ and the state derivatives $\dot{x}$. These correlations depend on the data, the system under study and chosen library of basis functions. Many, highly correlated terms can complicate the identification procedure with many models able to produce predictions close to the observed data. In contrast, in the case where only a few terms have high correlation, the identification of models with few terms is simplified. This effect could also explain the parsimonious models in the first two case studies.
@@
The choice of prior over models $p(m)$ may be critical in enforcing parsimony. Here, a strongly regularising geometric distribution parameterised by the number of terms in the model has been applied successfully. However, vastly many more parametrisations are possible. It is expected that prior selection in practice will be driven by expert domain knowledge, wherever possible. For example, one could set the prior over each term directly, assigning high prior probability to terms that are expected to be in the model and low probabilities otherwise. Alternatively one could promote parsimony in some types of library functions while permitting dense models in other types. The flexible nature of the Bayesian formulation presented in this work, allows modellers to encode both domain knowledge and cognitive biases a priori. An interesting avenue for further investigation is hierarchical prior structures whereby the model-space prior is permitted to be conditional on the values its parametrisation.
@@ Appendix C Robustness of BINDy
Nevertheless, it is instructive to consider the effect of low-data and high-noise regimes. In Figures 10 and 11, the posterior term inclusion probabilities are plotted. BINDy is applied to the Lorenz case-study example with a flat model prior. In Figure 10 the length of the observed data is held constant at 10s (or 1000 observations) and the variance of the added noise is varied from 1% to 15%. As can be seen in the Figure, BINDy is very robust to noise and assigns high posterior probability to correct terms in the data-generating model, with some spurious and missed terms assigned low probability at the highest noise levels.
Figure 10: Posterior term inclusion probabilities at a fixed data length for various levels of added Gaussian noise. Estimated by BINDy with a flat model prior. Subplot axes are identical to those of Figure 5. Solid bars indicate terms truly in the data generating model, while hollow bars indicate spurious terms.
In Figure 11, the effect of decreasing amounts of available data is investigated. The level of added noise is held constant at 2.5% RMS and the length of the observed dynamics available for inference is varied between 1.0 seconds (corresponding to only 100 datapoints) up to 10 seconds (corresponding to 1000 datapoints). The proposed BINDy approach with a flat model prior is applied to the data. Once again BINDy performs excellently, correctly assigning high probability to terms in the Lorenz equations with some spurious and missed terms in the shortest data lengths.
Figure 11: Posterior term inclusion probabilities at a fixed noise level for various amounts of observed dynamics. Estimated by BINDy with a flat model prior. Subplot axes are identical to those of Figure 5. Solid bars indicate terms truly in the data generating model, while hollow bars indicate spurious terms.
Figure 12 depicts the robustness of BINDy to the effect of data length and noise in a heatmap format. Each element of the heatmap considers the probability of the true data-generating model under the posterior estimated by BINDy. As can be seen in the figure, the probability approaches unity for low noise and long sequences.
Figure 12: Heatmap depicting the posterior probability of the true model for different noise levels and data lengths.